# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 4.4 Summary
In summary, there are deep connections between causal emergence and some hot topics in artificial intelligence such as CRL, abstractions, and world models. On one hand, the machine learning framework for causal emergence identification has similar structures to the frameworks of CRL. And even the concepts can have one to one correspondence as shown in Table 1. Therefore, the ideas and techniques of both sides can learn from each other. On the other hand, the studies of causal emergence can provide new perspectives and insights for causal representations and world model based algorithms. For example, we can explain that what the representations and causal mechanisms learned are emergent causal laws. Maximization of EI may help to improve the efficiency of learning the causalities for both CRL and reinforcement learning agents. Finally, we found a reinforcement learning agent with the ability to learn causal emergence may exhibit the ability of actions for purpose and objectives.
5 Potential Applications in Complex Systems
In this section, we will explore how the problems discussed in the previous sections relate to our understanding of complex systems. On one hand, the methods of causal emergence identification and the associated AI algorithms can be employed in complex systems to discover emergent patterns and causal relationships. On the other hand, the concepts of causal emergence can aid in our comprehension of complex systems, particularly in addressing significant issues such as consciousness and free will. These ideas have the potential to provide valuable insights and approaches to tackle these complex and profound problems.
5.1 Potential Applications
Complex systems typically consist of a large number of interacting units, and these units and interactions form various hierarchical and self-similar structures in a complex manner Mandelbrot and Mandelbrot (1982). In particular, when we examine complex systems from different scales, various structures and patterns may emerge at different scales, and these patterns may further interact with each other. Thus, a complex network of causality may form, and the nodes in this network may not only include the individual units of the complex system but also contains the emergent patterns in distinct scales.
As a result, it is crucial to determine how we can identify and extract emergent patterns and causal relationships from vast amounts of data generated by complex systems.
5.1.1 Emergence detection
In complex dynamical systems, various higher-level patterns, such as waves, periodic oscillations, and solitons, are ubiquitous. For instance, in the climate system, typhoons and tornadoes are emergent vortex structures. In urban areas, traffic jams also emerge as a result of interactions among a large number of cars. Social riots are another example of emergent events in higher levels of human society. Identifying these emergent patterns at an early stage is crucial and significant Haugen et al. (2023); Mnif and Müller-Schloer (2011). Therefore, there is an urgent need for a method that can automatically detect these emergent patterns and even provide early warning signals.
In O’toole et al. (2017), the authors regard emergent phenomenon as a specific event happened in a complex system, and detecting emergence is equivalent to detecting the formation and evaporation of this event. Therefore, emergence detection problem can be converted as a change-point detection problem of time series data.
Two methods are discriminated to detect emergence: one is centralized method, the other is decentralized. For example, in Mnif and Müller-Schloer (2011), the authors propose that the negative Shannon entropy can be used to measure the orderness of the entire system, and it can also indicate if emergence occurs. NIS and NIS+ are also centralized methods.
Another one is decentralized emergence detection method, in which the detectors are dispersed on the complex system, and implement event detection spontaneously and separately O’Toole et al. (2014).
As the development of artificial intelligence, some deep learning methods can be used to detect emergence. For example, dynamical image segmentation can be used to find interesting patterns in the video automatically Liu et al. (2021). A complex system can always be represented as a dynamical graph, e.g., a flock of birds, and when emergence occurs, the connectivity of this interaction network always increase. Therefore, we can detect the change of the network to find emergence Pazho et al. (2022). Other deep learning methods could be helpful.
5.1.2 Discovering emergent causality
Complex systems can be understood as a large network with events connected by causal links. A small disturb of one event may give rise a large roam of changes as butterfly effect indicates. While, sometimes, one single event happens with destiny which means it may be the effect of many different causes. How to discover the complex causal relationships from data in an automatic way is a challenging problem.
The research area of causal discovery tries to tackle this problem in many different ways Glymour et al. (2019); Zanga et al. (2022). For example, some studies try to use well-defined statistical indicators to characterize the causality between two variables, e.g., the measures of mutual information Donges et al. (2009), transfer entropy Stetter et al. (2012), etc. Nevertheless, these methods can not distinguish causal relationship from correlations. Another widely used method is independent causal discovery method Glymour et al. (2019), which tries to find the conditional independence among three variables in data, and derive Markov equivalent sub-graphs accordingly. Finally, binary causal interactions can be identified from these sub-graphs Kalisch and Bühlman (2007). Score based method tries to maximize or minimize a score function to find the optimal causal structures Zheng et al. (2018); Zhu et al. (2019). These methods always require a large amount of data and computation.
However, more challenges will be confronted when these causal discovery methods are applied in complex systems because circular causal structures and cross-level causality exist.
@@ 5.1.2 Discovering emergent causality
Circular causal structures widely exist in complex dynamical systems due to the feedback effect. That is, one variable may take effect on itself via a direct self-loop feedback or a circle with long chain. This circular causal structure may challenge the existed study methods of causality because most of these methods study directed acyclic causal graphs. Recent progress has developed some methods to discover these circular structures e.g.  Richardson (2013); Lacerda et al. (2012); Forré and Mooij (2018); Rantanen et al. (2020); Zhang et al. (2019); Pamfil et al. (2020); Vowels et al. (2022).
@@
Another aspect is higher level or cross-level causation if the ingredients of scale and coarse-graining are considered. For example, downward causation describes the causal effect between macro-level and micro-level variables. Thus, the causal connections may be cross-leveled. If causal emergence occurs in a complex system, some strong causality may be found between macro-variables. All of these factors in complex systems present new challenges. The machine learning methods discussed in section 4.1.2 can address the question of whether there exists higher-level causality at the macro-level. However, it’s important to note that this is a global property. It must be further developed and extended to find local causal relationship between macro-variables or macro-variable and micro-variable Varley (2022). While, existed methods of causal discovery must be extended to take the method of grouping a set of variables or coarse-graining a system into consideration. For example, the article Cheng et al. (2023) proposes a coarse-to-fine causal discovery algorithm based on Granger causality and graph neural network. The grouping process of variables is from coarse-to-find which can improve the efficiency of the algorithm. Other multi-level causal discovery methods are also proposed and discussed in Wang et al. (2023); Zhang et al. (2023); Fan et al. (2023).
One of the potential applications of these complicated causal discovery method is to enhance the interpretabiity of large and complex neural networks. As we know, interpretability is one of the notorious problem for large neural networks, although they are very useful to solve many real problems. Thus, we can apply the methods of causal discovery, especially, the methods to discover emergent causality on these neural network models to find hidden and emergent causal relationships between neurons or group of neurons Chattopadhyay et al. (2019); Ren et al. (2023). If the degrees of causal effect can be computed and compared among these discovered causal relationships, the most important causation can be pointed out to solve a concrete problem. This may enhance the interpretability of the neural network. In this way, the misery of emergent ability of large language models may be dismissed by uncovering the hidden causation within large models Michaud et al. (2023); Wei et al. (2022).
5.1.3 Beyond causality
In this review, we mainly focus on the emergence of causality, however, we do not need to be bounded on this kind of special emergence, more macroscopic or global level properties could also be emergent although causal emergence may be the most important one. For example, individuality is one of the emergent property with no any connection with causality which has been discussed in previous section. In  Chang et al. (2020), the authors discussed the concept of emergent information closure, which refers to the idea that the encoded information within an agent can form a closed system that is not influenced by the external world. The authors then argued that this agent with information closure be regarded as a model of consciousness.
Symmetry may be another interesting property within some complex systems, and it may be emergent. For example, any single trajectory of a large number of random walkers in a two dimensional Euclidean space is random, nevertheless, the Gaussian distribution surface describing how the number of walkers falling in each small region changes with different locations is isotropic. This kind of symmetry can only be found on the macro-level, thus, is emergent.
It should be noted that EI can also be viewed as a unique indicator of symmetry, specifically, it is time reversal symmetry. The reason behind this is that when the macro-dynamics approaches a time reversible map (a bijection between states across a time step), EI is increased. And the maximum EI can be arrived when the Markov dynamics is reversible. This interesting fact should be further studied in the future.
It is interesting and useful to deploy a machine learning method to automatically discover this kind of emergent symmetry. The framework of NIS may be extended to account for the problem of emergent symmetry identification if the optimized objective, EI, is replaced with some measure of isotope or more general symmetry Liu and Tegmark (2022). Similarly, we can also find other emergent symmetry with this framework once the symmetry can be measured and optimized.
5.1.4 Designing Emergence
We even can design a set of interaction rules for a complex system such that the desired property can emerge from the interactions once the desired property can be defined in a clear and quantitative way. The basic idea to realize emergence design is by optimization Adamatzky et al. ; Kreyssig and Dittrich (2011). Actually, the early studies of evolutionary computation or genetic algorithm have solved this problem for some simple and small cases Adamatzky et al. . For example, M. Mitchell et al. used genetic algorithm to search the interaction rules of one dimensional cellular automata such that the automata can solve some simple mathematical problems Mitchell et al. (1993). Recently, as the development of machine learning and automatic differential, searching such interaction rules may become easier and more efficient if the rules can be converted into differentiable operations Mordvintsev et al. (2022); Gala et al. (2023); Zhang and Walker (2022).
However, if the space of possible interaction rules is very large, all the methods based on search or optimization may be invalid. Thus, some ideas developed in the field of deep learning can be borrowed here to design emergence. For example, the framework of generative adversarial network (GAN) can be used to design complex systems to realize some emergent properties Goodfellow et al. (2014). On one side, we should design a generative model for a complex system or the interaction rules for the system such that a large number of candidate rules or systems can be generated. On the other side, a discriminator or a critic can be designed to evaluate the candidates if the desired emergent properties have been realized. Similar ideas can be used on other possible generative models, e.g., diffusion model, normalizing flow Cao et al. (2022); Yang et al. (2022); Papamakarios et al. (2021); Kobyzev et al. (2020), etc.
However, the difficulty of these methods lies in the demand for a large amount of training data and computations. The lack of data may hamper the applications. In general, designing emergence is a useful problem but it is also difficult.
@@ 5.2 Understanding Complex Systems from Causal Emergence
Deep understanding of causal emergence and emergent causality will help us to understand the hidden mechanism of complex systems including living systems, social systems, climate systems, ecosystems, etc. Particularly, some big problems like free will, consciousness Tononi and Koch (2015), and life are all related to emergent causation Walker (2014). For example, free will can be understood as an emergent downward causation Yurchenko (2023). In Chinese medicine, Yin and Yang can be understood as macro-states and the properties of their interactions are emergent causality. Social phenomena can be understood from the viewpoint of causal emergence Elder-Vass (2010). For instance, the increase of the rate of suicide may be the emergent causal effect of economic crises. Interestingly, both EI and $\phi$ID frameworks on causal emergence have relation with one of the consciousness theory: integrated information theory Tononi et al. (2016). As a result, further studies of causal emergence and the connections with other big problems are required.
@@
For a particular system, there are three problems that should be addressed in future studies: 1. When does causal emergence occurs? 2. How emergent causality take functional effects on the system? 3. How emergent causality may change when the system is changed to adapt the environment? We will discuss these problems one by one in brief.
@@
We still do not know when will causal emergence occurs and how does the measure of causal emergence change with some key parameters of a system. In Yang et al. (2023), the authors have shown how the measure of causal emergence change with different noises in an extended Boid model. It increases as the observational noise increases because coarse-graining operation can eliminate the observational noise. However, the same coarse-graining operation can hardly reduce the intrinsic noise (the noise within the dynamical equations), which leads to the degree of causal emergence decreases. It is reasonable to expect that there is a phase transition of the emergent causality within a complex system when some key parameter changes because the causality or the strength of causal effect is also a global property and it may dependent on some order parameters.
@@
If causal emergence occurs in the system, how does the emergent causation take effects on the part and the whole of the system? For example, the emergent "I" can be understood as an emergent macro-variable Yurchenko (2023). How does "I" can influence the part, say my foot, to move? This problem is non-trivial because it relates to the problem of mind-body interaction. One of the difficulty is that how can we build a model for such systems? Should we directly model the emergent properties and causation or model the underlying interaction rules and ignore the causation in higher level? If we directly model the emergent "I", the model may become very complex because the properties of "I" are difficult. However, if we model the system from the micro-level interactions, we can still hardly understand the emergent causality. Shall we switch the viewpoints between micro- and macro-? If yes, in what manner to switch?
@@
Finally, what is the relationship between adaptation and emergence Holland (1995, 2000, 1992)? Sometimes, when we refer to a property as emergent, we actually mean that this property can be developed through adaptation. Therefore, adaptation or evolution serves as a causal force for certain emergent properties. This concept also applies to emergent causality. For instance, downward causation, which is commonly found in complex systems, emerges as a result of adaptation and evolution. The next crucial question is: how can we evolve an emergent property or causation? This problem resembles the issue of designing emergence discussed in Section  5.1.4. However, our aim here is to seek an explanation rather than a design. We want to understand the specific environment and manner in which the observed emergent causation can be evolved through adaptation.
@@
All these problems desire further studies in future. However, these problems are just name a few, more interesting problems need to be addressed.
6 Conclusion
Quantifying the notion of emergence has been a challenge for a long time. Causal emergence serves as an important step in the development of a quantitative theory of emergence, making it a crucial concept in the field of complex system studies. In this paper, we provide a systematic review of papers on causal emergence and other related quantitative studies pertaining to emergence and causality. Our analysis highlights two crucial problem areas: emergence quantification and causal emergence identification in data. We discovered intriguing connections between the latter problem and artificial intelligence, particularly in sub-fields such as causal representation learning and world-model based reinforcement learning. As artificial intelligence continues to advance, we anticipate that more puzzles related to causality and cross-level interactions within complex systems will be solved.
7 Appendix A - Calculation of EI with TPM

### 25 Feb 2024 v3

@@ 5.1.2 Differences
In summary, there are deep connections between causal emergence and CRL. On one hand, the machine learning framework for causal emergence identification exhibits similar structures to the frameworks for CRL. And even the concepts can establish one-to-one correspondence, as shown in Table 1. Therefore, the ideas and techniques of both sides can learn from each other.
On the other hand, studies of causal emergence can provide new perspectives and insights for causal representation learning algorithms. For example, we can interpret the learned representations and causal mechanisms as emergent causal laws. The maximization of EI may improve the efficiency of learning the causalities for both CRL and reinforcement learning agents.
5.2 Ontological and Epistemological Causality and Emergence
Although machine learning techniques have facilitated the learning of causal structures and models, as well as the exploration of emergent properties and causation, it is important to consider whether the results obtained through machine learning reflect ontological causality and emergence or if they are merely epistemological phenomena.
Here, ontological causality refers to the causal relationships that exist in objective reality, independent of our knowledge or understanding. Ontological causality explores the fundamental mechanisms and interactions that give rise to causal effects in the world. Similarly, ontological emergence is concerned with the objective existence of emergent properties and their underlying mechanisms.
Epistemological causality, on the other hand, focuses on our knowledge and understanding of causality. It deals with how we perceive, model, and explain causal relationships based on our observations and experiences, but the causality may not exist in the real world. Epistemological emergence, akin to epistemological causality, focuses on our comprehension and elucidation of emergent phenomena. In other words, epistemological causality and emergence depend on an observer, and different observers may have distinct perspectives about causality and emergence for a particular objective phenomenon.
There has been a longstanding debate regarding the ontological and epistemological aspects of causality and emergence throughout history Kim (1992); Crutchfield and Young (1989); Bonabeau and Dessalles (2011); Yurchenko (2023). The authors of Yurchenko (2023) highlighted that the concept of “causation” in the literature is often vague and should be differentiated into “cause” and “reason”, aligning with ontological and epistemological causality. “Cause” means genuine cause that sufficiently causes the effects (causal closure principle Clayton and Davies (2006) and exclusion principle Kim (2007)), whereas “reason” serves as a mere explanation for individuals to comprehend the effects. Reason may not possess the same level of rigor as genuine cause, but it does offer a certain degree of predictability that individuals find valuable in certain circumstances.
However, distinguishing between them, particularly when addressing specific phenomena, remains a challenging task. One such controversial concept is downward causation, which has sparked extensive discussions. The question of whether downward causation exists objectively or not remains open. In Rosas et al. (2020), Rosas argued that downward causation not only exists independently of particular observers but also provides a characterization of the phenomenon through quantified measures using information decomposition. However, Yurchenko proposed that it is important to differentiate between two separate concepts: causality and reasoning. According to this perspective, downward causation falls under the category of “reason” rather than causation. In this context, causality and reasoning can be seen as representing ontological and epistemological causality, respectively.
Similarly, debates persist regarding the nature of causal emergence. The question arises as to whether causal emergence is a genuine phenomenon that exists independently of particular observers Dewhurst (2021). In Section 3.5.4, we highlighted that different coarse-graining strategies can lead to distinct macro-dynamic mechanisms with varying measures of EI. In essence, different coarse-graining strategies can be seen as representing different observers. However, Hoel’s theory proposes a criterion to differentiate between coarse-graining methods, namely EI maximization. Consequently, for a given set of Markovian dynamics, only the coarse-graining strategy and corresponding macro-dynamics that maximize the EI measure can be considered as an objective outcome. Nevertheless, a challenge arises when multiple solutions exist that maximize EI, introducing a degree of subjectivity Dewhurst (2021). The authors of Zhang and Liu (2022); Yang et al. (2023) addressed this issue by introducing constraints on prediction errors in micro-states, which partially alleviates the problem. However, the question of whether optimal results have unique solutions remains an open problem that requires further investigation.
Similarly, although the incorporation of machine learning cannot resolve debates surrounding ontological and epistemological causality and emergence, it can provide objective standards that help mitigate subjectivity. This is because machine learning algorithms strive to optimize objective functions. Thus, machine learning agents can be viewed as “objective” observers that make judgments on causality and emergence. This represents an additional advantage of incorporating machine learning. However, the issue of unique solutions is crucial in this approach and warrants further attention.
Are the machine learning results ontological or epistemological? The answer is that the results are epistemological and depend on machine learning algorithms. However, this does not mean that all the results of machine learning are meaningless since if the learning agents are well trained and the defined mathematical objectives are effectively optimized, the results can also be considered objective and independent of the algorithms.
It is intriguing to pose the same question regarding human cognition. Do all cognitive results in the human brain, encompassing causality, coarse-graining strategies, and macro-dynamics, reflect objective reality, or are they purely epistemological? Firstly, answering this question is exceedingly challenging. Secondly, the cognitive results specific to a particular human brain must objectively manifest within the neural network structures. Thus, for that brain, these results are ontological. Similar reasons hold for machine learning. The cognitive results of machine learning algorithms are recorded by the structures and parameters in neural networks. Therefore, studying these structures may reveal the properties of the original systems, and the nature of the interactions between the observer and the observed objects can be objectively reflected by machines.
Furthermore, the integration of machine learning can assist in establishing a theoretical framework for modeling observers and studying the interactions between the observer and the corresponding observed complex systems. This framework allows us not only to explore the hard problems regarding causality and emergence but also to understand the limitations and boundaries of the observer. One example of this research is the emergent classicality of quantum systems from the information bottleneck formed by machine learning algorithms Zhang and You (2023).
5.3 Potential Applications in Complex Systems
In Section 4, we discussed the techniques related to causal emergence identification, machine learning, and causal inference. In this sub-section, we discuss the potential applications of these techniques in complex systems. Complex systems can be understood as a large network, with events connected by causal links. Automatically discovering the complex causal relationships from data is a challenging problem.
The research area of causal discovery has tried to address this problem in various ways Glymour et al. (2019); Zanga et al. (2022); Donges et al. (2009); Kalisch and Bühlman (2007); Zheng et al. (2018); Zhu et al. (2019). However, additional challenges may arise when applying these causal discovery methods in complex systems because circular causal structures and cross-level causation may exist.
@@ 5.3 Potential Applications in Complex Systems
Circular causal structures widely exist in complex dynamical systems due to the feedback effect. That is, one variable may affect itself via a direct self-loop feedback or a long chain forming a circle. This circular causal structure may pose a challenge to existing methods of causality because most of these methods study directed acyclic causal graphs. However, recent progress has seen the development of some methods for discovering these circular structures in a data-driven manner, e.g., Richardson (2013); Lacerda et al. (2012); Forré and Mooij (2018); Rantanen et al. (2020); Zhang et al. (2019); Pamfil et al. (2020); Vowels et al. (2022); Spirtes (2013).
@@
Another aspect is that higher-level or cross-level causation may exist if the ingredients of scale and coarse-graining are considered. For example, downward causation describes the causal effect between macro-level and micro-level variables. Thus, the causal connections may be cross-level. If causal emergence occurs in a complex system, some strong causality may be found between macro-variables. All of these factors in complex systems present new challenges. The machine learning methods discussed in Section 4.1.1 can address the question of whether higher-level causality exists at the macro-level. However, it is important to note that this is a global property. It must be further developed and extended to find a local causal relationship between macro-variables or macro-variables and micro-variables Varley (2022). Moreover, existing methods of causal discovery must be extended to consider the method of grouping a set of variables or coarse-graining a system. For example, Cheng et al. (2023) proposed a coarse-to-fine causal discovery algorithm based on Granger causality and a graph neural network. The grouping process of variables is from coarse to fine, which can improve the efficiency of the algorithm. Other multi-level causal discovery methods were proposed and discussed in Wang et al. (2023); Zhang et al. (2023); Fan et al. (2023).
Another interesting problem is emergence detection. In complex dynamical systems, various higher-level patterns, such as waves, periodic oscillations, and solitons, are ubiquitous. For instance, in the climate system, typhoons and tornadoes are emergent vortex structures. In urban areas, traffic jams also emerge as a result of interactions among a large number of cars. Social riots are another example of emergent events in higher levels of human society. Identifying these emergent patterns at an early stage is crucial and significant Haugen et al. (2023); Mnif and Müller-Schloer (2011); O’toole et al. (2017); Mnif and Müller-Schloer (2011); O’Toole et al. (2014); Pazho et al. (2022); Liu et al. (2021). Therefore, there is an urgent need for a method that can automatically detect these emergent patterns and even provide early warning signals.
In this review, we mainly focus on the emergence of causality; however, we are not limited to this specific type of emergence. More macroscopic or global-level properties could also be emergent, although causal emergence may be the most important one. For example, in Chang et al. (2020), the authors discussed the concept of emergent information closure, which refers to the idea that the encoded information within an agent can form a closed system that is not influenced by the external world. The authors then argued that this agent with information closure can be regarded as a kind of consciousness. Symmetry may be another interesting property within some complex systems, and it may be emergent. For example, any single trajectory of a large number of random walkers in a two-dimensional Euclidean space is random. Nevertheless, the Gaussian distribution surface describing how the number of walkers falling in each small region changes with different locations is isotropic. This kind of symmetry can only be found at the macro-level and, thus, it is emergent. It would be interesting and useful to deploy a method for automatically discovering this kind of emergent symmetry. The framework for the NIS can be extended to address the problem of emergent symmetry identification if the optimized objective, EI, is replaced with a measure of isotope or more general symmetry Liu and Tegmark (2022). Similarly, we can also find other emergent symmetries using this framework once their measurement and optimization are possible.
However, it is crucial to acknowledge the inherent limitations of applying machine learning techniques to causal discovery and emergence identification problems. For instance, in Textor et al. (2016), the authors highlighted the existence of ’statistically equivalent’ but causally distinct DAGs, implying that different causal structures can be constructed to account for the same dataset. Consequently, when utilizing machine learning techniques to uncover causality and emergent properties, similar challenges need to be addressed. Further research in this direction warrants significant attention to overcome these obstacles.
@@ 5.4 Understanding Complex Systems from Causal Emergence
A profound understanding of causal emergence and emergent causality can provide insights into understanding various mysterious phenomena in complex systems, including living systems, social systems, climate systems, ecosystems, and more. Significantly, fundamental questions, such as free will Yurchenko (2023), consciousness Tononi and Koch (2015), and life, are all intricately connected to emergent causation Walker (2014). For instance, free will can be viewed as an emergent form of downward causation Yurchenko (2023). Social phenomena can be comprehended through the lens of causal emergence Elder-Vass (2010). Interestingly, both the EI and $\phi$ID frameworks for causal emergence have connections with one of the theories of consciousness, namely the integrated information theory Tononi et al. (2016). However, understanding these abstract concepts and phenomena in the context of causal emergence is merely an initial step.
@@
For a specific system, there are three problems that should be addressed in future studies: (1) When does causal emergence occur? (2) How does emergent causality have functional effects on the system? (3) How does emergent causality change when the system is changed to adapt to the environment? We discuss these problems one by one.
@@
We still do not know when causal emergence will occur and how the measure of causal emergence changes with some key parameters of a system. In Yang et al. (2023), the authors showed how the measure of causal emergence changes with different noises in the Boid model (see the relevant discussion in Section 4.1.2). It is reasonable to expect that there is a phase transition of emergent causality within a complex system when some key parameter changes because the causality or strength of the causal effect is also a global property, and it may be dependent on some order parameters.
@@
If causal emergence occurs in the system, how does the emergent causation affect the parts and the whole of the system? For example, the emergent “I” can be understood as an emergent macro-variable Yurchenko (2023). How does “I” influence the part, say the foot, to move? This problem is non-trivial because it relates to the problem of mind–body interaction. It is important to study the information flows at both the macro-level and micro-level together to understand this phenomenon.
@@
Finally, what is the relationship between adaptation and emergence Holland (1995, 2000, 1992)? Sometimes, when we refer to a property as emergent, we essentially mean that this property can be developed through adaptation. Therefore, adaptation or evolution serves as a causal force for certain emergent properties. This concept also applies to emergent causality. For instance, downward causation, which is commonly found in complex systems, emerges as a result of adaptation and evolution. The next crucial question is, how can we evolve an emergent property or causation? This problem resembles the issue of designing emergence Adamatzky et al. ; Kreyssig and Dittrich (2011); Mitchell et al. (1993). However, our aim here is to seek an explanation rather than a design. We want to understand the specific environment and manner in which the observed emergent causation can be evolved through adaptation.
@@
All these problems require further studies in the future. However, these problems are just a few of the more interesting problems that need to be addressed.
[…]
Conceptualization, J.Z.; methodology, J.Z., B.Y., P,C,; writing, B.Y., J.Z., A.L., J.W., Z.W., M.Y., K.L., M.M. All authors have read and agreed to the published version of the manuscript.
[…]
This work was supported in part by National Natural Science Foundation of China (No. 62141607, U1936219).
[…]
Not applicable.
[…]
No data available.
Acknowledgements.
[…]
Appendix A Calculation of EI with Transition Probability Matrix